OpenStudy (mitu12):

what is the prime factorization for the monomial without using exponents: 18xy

OpenStudy (mathstudent55):

What are the prime factors of 18?

OpenStudy (mitu12):

2*2*3*y*x

OpenStudy (anonymous):

HI mitu12. I believe you were the one that I had helped earlier? Factoring is really simple if you understand the basic idea that all numbers, (composite) are built up with prime numbers. Like 36 is made up of 2 x 2 x 3 x 3. It's just a matter of splitting it up. So, lets take a look at your problem. 18 can be divided by 2 = 9 9 can be divided by 3 = 3 2 x 3 x 3 = 18 "X" and "Y" are pretty much prime numbers themselves basically because we cannot factor them any more than we can factor the number 2. Your answer should be 2 x 3 x 3 x "x" x "y" Hopefully this helps.

OpenStudy (mitu12):

yes

OpenStudy (mathstudent55):

2 * 2 * 3 = 12 2 * 3 * 3 = 18 = 2 * 3^2 The prime factorization of 18xy is 2 * 3^2 * x * y